You are testing a new amusement park roller coaster with an empty car with a mass of 110 kg

kg . One part of the track is a vertical loop with a radius of 12.0 m

m . At the bottom of the loop (point A

A) the car has a speed of 25.0 m/s

m/s and at the top of the loop (point B

B) it has speed of 8.00 m/s

m/s .


1
Expert's answer
2022-03-21T12:46:34-0400

Answer

Force At the point A

"F=\frac{mv^2}{r}\\=\frac{110*25*25}{12}\\=5729.17N"

Force at B

"F'=\frac{mv^2}{r}\\=\frac{110*8*8}{12}\\=586.67N"